### v1

@@ 4.2.1 Datasets and Models
ADAS. Concerning advanced driving assistance, the DL system controls a vehicle in the Udacity simulator udacity-simulator , a cross-platform driving simulator developed with Unity3D unity , widely used in the ADS testing literature 2020-Stocco-ICSE ; 2020-Riccio-EMSE ; 2021-Jahangirova-ICST ; 10.1007/s10515-021-$0-$ The DL system includes a DAVE-2 model (a DNN regressor), a Level 2 iso26262 ADAS that performs the lane keeping functionality from a training set of images collected when the driver is an expert human pilot, by predicting the corresponding driving commands imitating the human driving behavior. We obtained the trained DAVE-2 model and the simulator from the replication package of existing work 2024-Biagiola-EMSE . The model architecture includes three convolutional layers for feature extraction, followed by five fully connected layers. The simulator supports the creation of open-loop road tracks for testing ADAS models, including the ability to generate and modify road topologies.
4.2.2 Comparison Baseline
To assess the guidance provided by our XAI-based approach, we compare XMutant against DeepJanus, a popular semantic-based test generator 2020-Riccio-FSE . For the selection of semantic concepts and mutation directions, DeepJanus relies on a random method that selects a candidate semantic concept for mutation based on a uniformly weighted probability. The mutation direction is task-specific. For IMDB, it applies one of the mutation methods on a randomly selected semantic concept. For MNIST, the mutation direction is sampled from the uniform distribution over 00 to $2\pi$. For ADAS, the mutation direction is randomly selected from one of the two directions perpendicular to the road’s curvature. To ensure a fair comparison, we applied the same extent value for mutation direction to both XMutant and DeepJanus’s mutation methods. In our study, we sampled the extent of MNIST from a uniform distribution over the interval $[0,1.2]$, and for ADAS we adopted a fixed value 4, proportional to the road’s width 8. These values are determined based on existing literature 2020-Riccio-FSE ; 2024-Biagiola-EMSE .
@@ 4.2.3 Metrics used for Analysis
Concerning RQ1, we evaluate the effectiveness of the test generation technique by computing the cumulative failure rate observed under a given number of mutation iterations across different configurations. The cumulative failure rate is obtained by dividing the number of failures by the total number of inputs. For IMDB and MNIST, a failure is characterized by the number of misclassifications. For ADAS, we measure the number of safety-critical failures in terms of out of bounds (OOBs), which occur when the vehicle drives outside the road’s drivable lanes during the execution in the simulator.
Concerning RQ2, we evaluate the performance of XMutant by measuring the relative efficiency with respect to DeepJanus. Particularly, relative efficiency is computed by integrating the cumulative failure rates over the number of iterations to obtain the area under the cumulative failure curve (AUFC). For each configuration and number of iterations, we calculate the AUFC (Area Under the Failure Curve) and then divide it by the AUFC of our comparison baseline. Considering computational time complexity, XMutant introduces an additional overhead for computing local explanations compared to the baseline. To assess its impact, we specifically evaluated the proportion of the XAI overhead relative to the baseline’s average computation time per iteration.
@@ 4.2.4 Configurations
We evaluate a total of 24 configurations, varying the XAI algorithm, the semantic concept selection method, and the mutation direction as described next.
@@
In the MNIST classification task, XMutant uses three XAI algorithms (SmoothGrad smilkov2017smoothgrad , Grad-CAM++ Grad-CAM++_2018 , and Integrated Gradients integratedgradients ) and two strategies (square windows and clustering). For square windows, we use a window size value of $\textit{ws}=3$, which was found appropriate during preliminary experiments (e.g., such value does not lead to the creation of square windows that exceed the input boundaries). For mutation direction, XMutant relies on random selection due to the absence of attractor information. The clustering strategy allows for the mutation direction to be chosen either towards the attention attractor or randomly. Additionally, XMutant pre-processes the heatmap to discard low-intensity values and filter out the noise using a threshold $\epsilon=0.1$, which was found adequate from preliminary experiments.
@@ 4.3 Results
4.3.1 Effectiveness, efficiency, configuration (RQ1, RQ2 and RQ3)
Table 2 presents the effectiveness, performance, and comparative results for all configurations of XMutant and DeepJanus as the baseline, for all case studies (IMDB, MNIST, and ADAS). The table reports the information at different iteration intervals, respectively 10 for IMDB, 100 for MNIST, and 5 for ADAS. For each configuration of XMutant and DeepJanus, the first number denotes the cumulative failure rate, whereas the second number indicates the relative efficiency with respect to the baseline.
@@ 4.3.1 Effectiveness, efficiency, configuration (RQ1, RQ2 and RQ3)
Concerning effectiveness (RQ1), for all case studies, all configurations of XMutant outperform the baseline, regardless of the iteration considered. We assessed the statistical significance of the differences in the cumulative failure rate between XMutant and the baseline using the non-parametric Mann-Whitney U test Wilcoxon1945 (with $\alpha=0.01$) and the magnitude of the differences using the Cohen’s $d$ effect size cohen1988statistical . The differences were found to be statistically significant ($p$-value $<$ 0.01), with varying effect sizes.
@@
Table 2: Results for RQ1, RQ2 and RQ3. For each configuration, the two values indicate the Cumulative Failure Rate [%]/ Relative Efficiency over different Iterations. P-values indicating statistically significant differences are highlighted in bold.
@@
RQ1: The guidance provided by the XAI allows XMutant to generate significantly more failure-inducing inputs than DeepJanus (up to $+208\%$ for sentiment analysis, up to $+125\%$ for digit recognition, and $+27\%$ for system-level advanced driving assistance).
Concerning performance (RQ2), all configurations of XMutant were producing misbehaviors faster than the baseline, as evidenced by the relative efficiencies regarding DeepJanus as the relative efficiencies are greater than 1. The results hold for all case studies. We also perform the statistical tests for RQ2 on the iteration counts between XMutant and the baseline, which revealed that in 17 out of 21 configurations of XMutant, the differences are statistically significant ($p$-value $<$ 0.01) compared to DeepJanus.
Figure 4 illustrates the time required to compute the local explanations as a percentage of the baseline’s average computation time per iteration. Notably, LIME is extremely time-consuming since it requires estimating a surrogate model, making it unsuitable for test generation. Aside from LIME, XMutant with other XAI methods remains more efficient than DeepJanus, even when the overhead is accounted for. Besides, in system-level ADAS testing, XAI computations constitute a relatively small portion of the total execution time, as executing the driving simulation is computationally expensive. These results show XMutant’s time-efficiency also for complex system-level testing settings.
RQ2: XMutant is faster than DeepJanus at exposing failure-inducing inputs for DL systems (ranging from 2$\times$ to 7$\times$ times faster for model-level sentiment analysis and digit recognition, and 2$\times$ faster for system-level advanced driving assistance).
Figure 4: Computational overhead for local explanation.
4.3.2 Validity (RQ4)
(a) IMDB.
(b) MNIST.
4.3.3 RQ5 (Comparison)
Technique Effectiveness Efficiency (s) Density Coverage
DLFuzz 60.10% 0.65 0.2156 0.336
FGSM 43.80% 0.37 0.0104 0.05
DeepXplore 0.02% 76.84 0.2591 0.11
XMutant 74.15% 0.33 0.3824 0.41
4.4 Threats to Validity
4.4.1 Internal validity
4.4.2 External validity
5 Qualitative Analysis
(a) Qualitative analysis for MNIST by SmoothGrad
(b) Qualitative analysis for MNIST Grad-CAM++
5.1 Qualitative Analysis of Digit Recognition
We analyzed two XAI methods, SmoothGrad (7(a)) and Grad-CAM++ (7(b)), selecting three mutation configurations guided by the same seed. The first observation is related to gradient attenuation, as the mean intensity of both methods decreases with the confidence, especially those driven by XMutant. Notably, Grad-CAM++’s mean intensity drops significantly, thus it is visualized in a logarithmic scale. Since SmoothGrad reflects gradients across the entire model, the gradient attenuation is less pronounced w.r.t. Grad-CAM++, which instead focuses on local gradients at the last convolutional layer, directly exposing the gradient attenuation in this layer and resulting in a more pronounced decline. Our second observation is related to entropy degradation. The inputs generated by XMutant cause the entropy to increase when the confidence decreases, indicating that the DNN progressively loses its focus. This phenomenon is also more observable with the mutation guided by XMutant equipped with SmoothGrad, since it generates noise-averaged gradients, whereas the entropy of heatmap generated by Grad-CAM++ appears noisy.
5.2 Qualitative Analysis of ADAS
Figure 8 depicts the distribution of DQD, i.e., the subtraction of driving quality pairs. DeepJanus produces only minor variations, with smaller means and a tighter distribution, with its median closely aligned with symmetry around zero. On the other hand, the mutations led by XMutant result in larger mean and variance values for DQD, as well as more data points distributed in areas with high values, indicating a higher driving quality degradation due to more challenging road topologies. Within each approach, we conduct the above-mentioned statistical test over the driving quality distributions before and after mutation. The differences in the quality of driving before/after mutation were found to be statistically significant according to the non-parametric Mann-Whitney U test Wilcoxon1945 ($p$-value $<$ 0.01), which is not the case for the DeepJanus approach.
6 Discussion
Our approach depends on the quality of the local explanations, specifically on their capability to constitute a correct reference of the DL system’s behavior. Additionally, understanding the characteristics of these explanations by different XAI methods is crucial for XMutant to leverage them for effective mutation selection and direction. The results of this study showed the significant benefits of utilizing XAI for semantic-based fuzzing. XAI enables the identification of DNN’s misbehaviors while ensuring that the tests remain valid and relevant to the original input domain, thanks to the semantic representation and focused test generation. Although the generation of local explanation introduces additional computational overhead, our findings suggest that the trade-off is justified. Our effectiveness results, even for complex applications such as advanced driving assistance systems, indicate that the benefits of XAI in failure exposure outweigh the costs of computing heatmaps. Our experiments suggest that incorporating XAI (except for the model-agnostic XAI method LIME) into the testing workflow is beneficial. Additionally, the adopted open-source XAI methods can be further optimized for efficiency by parallelizing multiple backpropagations or DNN inferences. We will evaluate this aspect in our future work.
7 Related Work
7.1 Test Generation for Deep Learning Systems
Semantic Input Manipulation (SIM) techniques leverage a semantic representation of the input domain (e.g., a model) to generate test inputs, similar to conventional model-driven engineering practices that uphold compliance with domain-specific constraints Abdessalem-ASE18-1 ; Abdessalem-ASE18-2 ; Abdessalem-ICSE18 ; Gambi:2019:ATS:3293882.3330566 ; 2020-Riccio-FSE ; isa . The manipulation occurs on the model, which is subsequently reconverted to the original format larman1998applying . SIM techniques operate within a restricted input space, specifically the control parameters of the model representation. These techniques enhance the realism of the produced outputs by implementing appropriate model constraints.
Raw input manipulation (RIM) techniques involve modifying an image’s original pixel space to create a new input by perturbing the pixel values. RIM techniques aim to produce minimal, often imperceptible changes to original to trigger misbehaviors in the DL system. These methods target different aspects of testing, such as data augmentation or adversarial attacks, which are not directly aligned with our goal. Our method is a functional test generator, differing from adversarial testing in both goals and techniques. Functional testing creates new, valid, in-distribution inputs to evaluate a DNN’s generalization. In contrast, adversarial testing adds minor perturbations to original inputs to test robustness. However, for completeness, we describe the main propositions next.
Latent space manipulation (LIM) techniques generate new inputs by learning and reconstructing the underlying distribution of the input data. Sinvad kang2020sinvad constructs the input space using Variational Autoencoders (VAE) kingma2013auto ; doersch2016tutorial and navigates the latent space by adding a random value sampled from a normal distribution to a single element of the latent vector. Sinvad aims to explore the latent space by maximizing either the probability of misbehaviors, estimated from the softmax layer output, or by surprise coverage kim2019guiding . The Feature Perturbations technique dunn2021exposing ; DBLP:journals/corr/abs-2001-11055 involves injecting perturbations into the output of the generative model’s first layers, which represent high-level features of images. These perturbations can affect various characteristics of the image, such as shape, location, texture, or color. DeepRoad deeproad generates driving images using Generative Adversarial Networks (GANs) goodfellow2014generative ; goodfellow2020generative for image-to-image translation. CIT4DNN dola2024cit4dnn combines VAE and combinatorial testing cit605761 . This allows the systematic exploration and generation of diverse and infrequent input datasets. CIT4DNN partitions latent spaces to create test sets that contain a wide range of feature combinations and rare occurrences. A recently proposed technique, Instance Space Analysis, aims to pinpoint the critical features of test scenarios that impact the detection of unsafe behavior isa .
7.2 XAI for DL Testing
8 Conclusions
In this work, we describe and evaluate XMutant, a semantic-based fuzzer for DL systems that generates inputs that focus on the attention of the system under test through mutations that are informed by the local explanation available from XAI algorithms. We evaluated XMutant on both model-level and system-level testing. Our empirical studies show that XMutant is significantly more effective and efficient than the state-of-the-art test generation approaches while preserving a high validity rate of failing test inputs.

### v2

@@ 4.2.1 Datasets and Models
ADAS. Concerning advanced driving assistance, the DL system controls a vehicle in the Udacity simulator udacity-simulator, a cross-platform driving simulator developed with Unity3D unity, widely used in the ADS testing literature 2020-Stocco-ICSE; 2020-Riccio-EMSE; 2021-Jahangirova-ICST; 10.1007/s10515-021-00310-0. The DL system includes a DAVE-2 model (a DNN regressor), a Level 2 iso26262 ADAS that performs the lane keeping functionality from a training set of images collected when the driver is an expert human pilot, by predicting the corresponding driving commands imitating the human driving behavior. We obtained the trained DAVE-2 model and the simulator from the replication package of existing work 2024-Biagiola-EMSE. The model architecture includes three convolutional layers for feature extraction, followed by five fully connected layers. The simulator supports the creation of open-loop road tracks for testing ADAS models, including the ability to generate and modify road topologies.
4.2.2 Comparison Baselines
To assess the guidance provided by our XAI-based approach, we compare XMutant against DeepJanus, a popular semantic-based test generator 2020-Riccio-FSE. For the selection of semantic concepts and mutation directions, DeepJanus relies on a random method that selects a candidate semantic concept for mutation based on a uniformly weighted probability. The mutation direction is task-specific. For IMDB, it applies one of the mutation methods on a randomly selected semantic concept. For MNIST, the mutation direction is sampled from the uniform distribution over 0 to $2\pi$. For ADAS, the mutation direction is randomly selected from one of the two directions perpendicular to the road’s curvature. To ensure a fair comparison, we applied the same extent value for mutation direction to both XMutant and DeepJanus’s mutation methods. In our study, we sampled the extent of MNIST from a uniform distribution over the interval $[0,1.2]$, and for ADAS we adopted a fixed value 4, proportional to the road’s width 8. These values are determined based on existing literature 2020-Riccio-FSE; 2024-Biagiola-EMSE.
@@ 4.2.3 Metrics used for Analysis
Concerning RQ1, we evaluate the effectiveness of the test generation technique by computing the cumulative failure rate observed under a given number of mutation iterations across different configurations. The cumulative failure rate is obtained by dividing the number of failures by the total number of inputs. For IMDB and MNIST, a failure is characterized by the number of misclassifications. For ADAS, we measure the number of safety-critical failures in terms of out of bounds (OOBs), which occur when the vehicle drives outside the road’s drivable lanes during execution in the simulator.
Concerning RQ2, we evaluate the performance of XMutant by measuring the relative efficiency with respect to DeepJanus. Particularly, relative efficiency is computed by integrating the cumulative failure rates over the number of iterations to obtain the area under the cumulative failure curve (AUFC). For each configuration and number of iterations, we calculate the AUFC (Area Under the Failure Curve) and then divide it by the AUFC of our comparison baseline. Considering computational time complexity, XMutant introduces an additional overhead for computing local explanations compared to the baseline, while all other components in the fuzzing loop introduce the same computational expense. To assess its impact, we specifically evaluated the proportion of the XAI overhead relative to the baseline’s average computation time per iteration. To measure computational cost, we compute the XAI overhead, defined as the average time per iteration spent on local explanation, normalized by the baseline (DeepJanus) iteration time:
$\text{XAI overhead}=\frac{T_{\textsc{XMutant}}-T_{\text{baseline}}}{T_{\text{baseline}}}$
To jointly assess these two aspects, relative efficiency and computational cost, we also report a composite efficiency metric, which adjusts the relative AUFC by factoring in the per-iteration overhead, i.e.,
$\text{composite efficiency}=\frac{\text{relative efficiency}}{1+\text{XAI overhead}}$
offering a more operational view on real-world efficiency under time constraints.
@@ 4.2.4 Configurations
We evaluate a total of 24 configurations, designed to systematically assess the impact of the core components (XAI algorithm, the semantic concept selection method, and the mutation direction) of XMutant for the purpose of ablation-style analysis.
@@
In the MNIST classification task, XMutant uses three XAI algorithms (SmoothGrad smilkov2017smoothgrad, Grad-CAM++ Grad-CAM++_2018, and Integrated Gradients integratedgradients) and two strategies (square windows and clustering). For square windows, we use a window size value of $\textit{ws}=3$, which was found appropriate during preliminary experiments (e.g., such value does not lead to the creation of square windows that exceed the input boundaries), as presented in Appendix 10.1. For mutation direction, XMutant relies on random selection due to the absence of attractor information. The clustering strategy allows for the mutation direction to be chosen either towards the attention attractor or randomly, thereby isolating the effect of mutation direction. Additionally, following previous works fongInterpretableExplanations; selvaraju2017grad, XMutant pre-processes the heatmap to discard low-intensity values and filter out the noise using a threshold $\epsilon=0.1$, improving stability of semantic selection.
@@
These controlled variations across tasks and configurations enable us to analyze the contribution of each component in guiding semantic-based fuzzing.
@@ 4.3 Results
4.3.1 Effectiveness (RQ1)
@@
Concerning effectiveness (RQ1), Table 2 presents the effectiveness result (cumulative failure rate) for all configurations of XMutant and DeepJanus as the baseline, for all case studies (IMDB, MNIST, and ADAS), over different iterations. For all case studies, all configurations of XMutant outperform the baseline, regardless of the iteration considered. We assessed the statistical significance of the differences in the cumulative failure rate between XMutant and the baseline using the non-parametric Mann-Whitney U test Wilcoxon1945 (with $\alpha=0.01$) and the magnitude of the differences using the Cohen’s $d$ effect size cohen1988statistical. The differences were found to be statistically significant ($p$-value << 0.01), with varying effect sizes.
@@
Table 2: Results for Effectiveness (RQ1). Values indicate the Cumulative Failure Rate [%] over different Iterations. P-values with statistically significant differences are boldfaced.
Iterations SmoothGrad LIME Integrated Gradients DJ
20 21.84 32.18 31.03 9.77
40 39.08 50.57 56.90 17.82
60 46.55 62.64 66.67 24.14
80 50.57 69.54 73.56 28.74
IMDB 100 54.60 76.44 75.86 32.18
p-value 6.93E-18 1.17E-17 7.90E-18 -
effect size large large large -
SmoothGrad Grad-CAM++ Integrated Gradients DJ
Clustering SW Clustering SW Clustering SW -
200 77 24 24 74 21 22 60 27 26 15
400 94 47 45 94 46 45 80 51 47 35
600 97 62 60 98 60 61 87 66 61 51
800 98 71 69 99 70 73 90 76 71 62
MNIST 1000 99 78 76 99 77 80 92 82 77 70
effect size large small small large small small large medium small -
SmoothGrad Grad-CAM++ Integrated Gradients DJ
Iterations Random Low High Random Low High Random Low High -
10 65 60 67 75 87 75 64 65 58 51
20 89 82 84 95 93 95 87 85 85 75
30 91 89 95 95 95 95 95 87 89 84
ADAS 40 95 89 95 95 95 95 95 93 95 95
effect size small small small medium medium medium small small small -
Table 3: Results for Efficiency (RQ2). Values indicate the Relative Efficiency (AUFC Ratio) over different Iterations, followed by XAI Computation Overhead and composite time-to-failure efficiency compared to the DeepJanus. The Iteration used for the computation of composite efficiency is underlined. The composite efficiency values outperforming baseline are boldfaced.
Iterations SmoothGrad LIME Integrated Gradients DJ
10 1.93 1.44 2.05 1
20 2.08 2.31 2.71 1
40 2.12 2.82 3.05 1
IMDB 100 1.90 2.74 2.59 1
XAI overhead $\downarrow$ 0.35 236.63 1.47 0
Composite Efficiency $\uparrow$ 1.57 0.01 1.23 1
SmoothGrad Grad-CAM++ Integrated Gradients DJ
Clustering SW Clustering SW Clustering SW -
100 7.00 1.83 2.11 4.81 1.03 1.32 5.54 2.06 2.15 1
200 6.37 1.77 1.82 5.54 1.32 1.48 4.92 2.00 1.97 1
400 4.03 1.52 1.47 3.84 1.34 1.36 3.26 1.67 1.57 1
MNIST 1000 2.12 1.23 1.20 2.09 1.19 1.21 1.86 1.32 1.23 1
XAI overhead $\downarrow$ 0.87 0.87 0.87 0.79 0.79 0.79 0.52 0.52 0.52 0
Composite Efficiency $\uparrow$ 3.74 0.98 1.13 2.69 0.58 0.74 3.64 1.36 1.41 1
SmoothGrad Grad-CAM++ Integrated Gradients DJ
Iterations Random Low High Random Low High Random Low High -
5 2.00 2.20 2.87 2.80 2.53 2.07 2.13 2.53 2.60 1
10 1.35 1.36 1.69 1.86 1.78 1.60 1.57 1.60 1.56 1
20 1.31 1.21 1.36 1.47 1.53 1.48 1.30 1.30 1.32 1
ADAS 40 1.15 1.09 1.18 1.23 1.25 1.23 1.16 1.12 1.14 1
XAI overhead $\downarrow$ 0.19 0.19 0.19 0.16 0.16 0.16 0.16 0.16 0.16 0
Composite Efficiency $\uparrow$ 1.68 1.85 2.41 2.41 2.18 1.78 1.83 2.18 2.24 1
RQ1: The guidance provided by the XAI allows XMutant to generate significantly more failure-inducing inputs than DeepJanus (up to $208\%$ for sentiment analysis, up to $125\%$ for digit recognition, and $27\%$ for system-level advanced driving assistance).
4.3.2 Efficiency (RQ2)
Concerning performance (RQ2), Table 3 presents (1) the relative efficiency, i.e., the ability to induce failures quickly, measured by the relative AUFC w.r.t. DeepJanus; (2) the XAI overhead, i.e., the additional per-iteration cost introduced by XAI computations as a ratio of the baseline’s average computation time per iteration; and the joint metric composite efficiency, which reflects the real-world time effectiveness by balancing AUFC gains and computational costs. Regardless of the XAI overhead, all configurations of XMutant were producing misbehaviors faster than the baseline, as evidenced by the relative efficiencies regarding DeepJanus as the relative efficiencies are greater than 1. The results hold for all case studies. Considering the XAI overhead, LIME is extremely time-consuming since it requires estimating a surrogate model, making it unsuitable for test generation. 17 out of 21 configurations of XMutant with other XAI methods remains more efficient than DeepJanus, even when the overhead is accounted for, reported by the composite efficiencies larger than 1. Besides, in system-level ADAS testing, XAI computations constitute a relatively small portion of the total execution time, as executing the driving simulation is computationally expensive. These results show XMutant’s time-efficiency, especially for complex system-level testing settings.
[…]
4.3.3 Configuration (RQ3)
4.3.4 Validity (RQ4)
(a) IMDB.
(b) MNIST.
4.3.5 RQ5 (Comparison)
Technique Effectiveness Efficiency (s) Density Coverage
DLFuzz 60.10% 0.65 0.2156 0.34
FGSM 43.80% 0.37 0.0104 0.05
DeepXplore 0.02% 76.84 0.2591 0.11
XMutant 74.15% 0.33 0.3824 0.41
Figure 6: An example of generated failure-inducing inputs by raw input manipulation techniques and XMutant.
Figure 6 compares the input produced by XMutant and the other three raw input manipulation methods (DLFuzz, FGSM, and DeepXplore). These pixel-level fuzzing approaches inject different patterns of noise, either in the global background or locally focused on some specific areas, while largely preserving the semantic content of the input (the shape of the digit) from the original seed. In contrast, XMutant operates on a semantic input representation, modifying the shape of the digit through control-point mutations. This leads to functionally new and visibly distinct test inputs challenging the model under test.
4.4 Threats to Validity
4.4.1 Internal validity
4.4.2 External validity
5 Qualitative Analysis
Figure 7: Qualitative analysis for MNIST by SmoothGrad (top) and Grad-CAM++ (bottom).
5.1 Qualitative Analysis of Digit Recognition
We analyzed two XAI methods, SmoothGrad and Grad-CAM++ (Figure 7), selecting three mutation configurations guided by the same seed. The first observation is related to gradient attenuation, as the mean intensity of both methods decreases with the confidence, especially those driven by XMutant. Notably, Grad-CAM++’s mean intensity drops significantly, thus it is visualized in a logarithmic scale. Since SmoothGrad reflects gradients across the entire model, the gradient attenuation is less pronounced w.r.t. Grad-CAM++, which instead focuses on local gradients at the last convolutional layer, directly exposing the gradient attenuation in this layer and resulting in a more pronounced decline. Our second observation is related to entropy degradation. The inputs generated by XMutant cause the entropy to increase when the confidence decreases, indicating that the DNN progressively loses its focus. This phenomenon is also more observable with the mutation guided by XMutant equipped with SmoothGrad, since it generates noise-averaged gradients, whereas the entropy of heatmap generated by Grad-CAM++ appears noisy.
5.2 Qualitative Analysis of ADAS
Figure 8 depicts the distribution of DQD, i.e., the subtraction of driving quality pairs. DeepJanus produces only minor variations, with smaller means and a tighter distribution, with its median closely aligned with symmetry around zero. On the other hand, the mutations led by XMutant result in larger mean and variance values for DQD, as well as more data points distributed in areas with high values, indicating a higher driving quality degradation due to more challenging road topologies. Within each approach, we conduct the above-mentioned statistical test over the driving quality distributions before and after mutation. The differences in the quality of driving before/after mutation were found to be statistically significant according to the non-parametric Mann-Whitney U test Wilcoxon1945 ($p$-value << 0.01), which is not the case for the DeepJanus approach.
6 Discussion
Although different XAI methods may vary significantly in their quality and behavior, particularly in terms of established evaluation metrics such as faithfulness and sensitivity li2021experimental; adebayo2018sanity, our results demonstrate that XMutant consistently improves failure discovery across most configurations. This suggests that XMutant does not require the explanations to be fully accurate or faithful, but rather benefits from the heuristic signal provided by these explanations to increase the likelihood of failure-inducing inputs. This is due to the nature of fuzzing itself, which is inherently exploratory and tolerant of imperfect guidance. Therefore, XMutant remains robust even when the underlying XAI techniques vary in their quality and behavior.
Additionally, understanding the characteristics of these explanations by different XAI methods is crucial for XMutant to leverage them for effective mutation selection and direction. For example, attribution-based methods are particularly useful for determining mutation direction, while sensitivity-based methods are helpful to guide selection toward more influential input components. XMutant adapts its mutation strategy accordingly to accommodate these differences in explanation semantics.
The results of this study showed the significant benefits of utilizing XAI for semantic-based fuzzing. XAI enables the identification of DNN’s misbehaviors while ensuring that the tests remain valid and relevant to the original input domain, thanks to the semantic representation and focused test generation. Although the generation of local explanation introduces additional computational overhead, our findings suggest that the trade-off is justified. Our effectiveness results, even for complex applications such as advanced driving assistance systems, indicate that the benefits of XAI in failure exposure outweigh the costs of computing heatmaps. Our experiments suggest that incorporating XAI (except for the model-agnostic XAI method LIME) into the testing workflow is beneficial. Additionally, the adopted open-source XAI methods can be further optimized for efficiency by parallelizing multiple backpropagations or DNN inferences. We will evaluate this aspect in our future work.
7 Related Work
7.1 Test Generation for Deep Learning Systems
Semantic Input Manipulation (SIM) techniques leverage a semantic representation of the input domain (e.g., a model) to generate test inputs, similar to conventional model-driven engineering practices that uphold compliance with domain-specific constraints Abdessalem-ASE18-1; Abdessalem-ASE18-2; Abdessalem-ICSE18; Gambi:2019:ATS:3293882.3330566; 2020-Riccio-FSE; isa; 2023-Stocco-EMSE; 2023-Stocco-TSE. The manipulation occurs on the model, which is subsequently reconverted to the original format larman1998applying. SIM techniques operate within a restricted input space, specifically the control parameters of the model representation. These techniques enhance the realism of the produced outputs by implementing appropriate model constraints.
Raw input manipulation (RIM) techniques involve modifying an image’s original pixel space to create a new input by perturbing the pixel values lambertenghi_ICST25. RIM techniques aim to produce minimal, often imperceptible changes to original to trigger misbehaviors in the DL system. These methods target different aspects of testing, such as data augmentation or adversarial attacks, which are not directly aligned with our goal. Our method is a functional test generator, differing from adversarial testing in both goals and techniques. Functional testing creates new, valid, in-distribution inputs to evaluate a DNN’s generalization. In contrast, adversarial testing adds minor perturbations to original inputs to test robustness. However, for completeness, we describe the main propositions next.
Latent space manipulation (LIM) techniques generate new inputs by learning and reconstructing the underlying distribution of the input data 2025-Guo-arxiv. Sinvad kang2020sinvad constructs the input space using Variational Autoencoders (VAE) kingma2013auto; doersch2016tutorial and navigates the latent space by adding a random value sampled from a normal distribution to a single element of the latent vector. Sinvad aims to explore the latent space by maximizing either the probability of misbehaviors, estimated from the softmax layer output, or by surprise coverage kim2019guiding. The Feature Perturbations technique dunn2021exposing; DBLP:journals/corr/abs-2001-11055 involves injecting perturbations into the output of the generative model’s first layers, which represent high-level features of images. These perturbations can affect various characteristics of the image, such as shape, location, texture, or color. DeepRoad deeproad generates driving images using Generative Adversarial Networks (GANs) goodfellow2014generative; goodfellow2020generative for image-to-image translation. CIT4DNN dola2024cit4dnn combines VAE and combinatorial testing cit605761. This allows the systematic exploration and generation of diverse and infrequent input datasets. CIT4DNN partitions latent spaces to create test sets that contain a wide range of feature combinations and rare occurrences. Instance Space Analysis, aims to pinpoint the critical features of test scenarios that impact the detection of unsafe behavior isa. Mimicry weissl2025targeted explores the behavioral boundaries of deep learning classifiers by generating boundary inputs through manipulation of disentangled latent representations learned by a style-based generative model.
7.2 XAI for DL Testing
8 Conclusions
In this work, we describe and evaluate XMutant, a semantic-based fuzzer for DL systems that generates inputs that focus on the attention of the system under test through mutations that are informed by the local explanation available from XAI algorithms. We evaluated XMutant on both model-level and system-level testing. Our empirical studies show that XMutant is significantly more effective and efficient than the state-of-the-art test generation approaches, while preserving a high validity rate of failing test inputs.
9 Data Availability
All our results, the source code, and the simulator are accessible and can be reproduced replication-package.
10 Appendix
10.1 Effect of Window Size
To validate the choice of the window size $ws=3$ for the XMutant equipped with Square-Window in MNIST study, we conducted an ablation study across different values (1, 3, 5, 7), as shown in Figure 9. We report the failure rates using both SmoothGrad and Grad-CAM++, for 2000 seeds and with the budget of 200 iterations.
Figure 9: Effect on Window Size
While SmoothGrad peaks around $ws=5$, Grad-CAM++ performs best at $ws=1$. On average, $ws=3$ yields the highest combined score across methods. However, the variation is relatively minor (within  2%), suggesting the method is not overly sensitive to this parameter. Based on this, we selected $ws=3$ for running our experiments to balance performance and generality.
10.2 Comprehensive result for RQ1, RQ2, and RQ3
Table 5: Detailed results for RQ1, RQ2 and RQ3. For each configuration, the two values indicate the Cumulative Failure Rate [%]/ Relative Efficiency over different Iterations. P-values indicating statistically significant differences are highlighted in bold.
@@ 10.2 Comprehensive result for RQ1, RQ2, and RQ3
Table 5 comprehensively presents the effectiveness, performance, and comparative results for all configurations of XMutant and DeepJanus as the baseline, for all case studies (IMDB, MNIST, and ADAS). The table reports the information at different iteration intervals, respectively, 10 for IMDB, 100 for MNIST, and 5 for ADAS. For each configuration of XMutant and DeepJanus, the first number denotes the cumulative failure rate, whereas the second number indicates the relative efficiency with respect to the baseline.
10.3 Sanity Check on LLM’s Validation
We conducted a sanity check of ChatGPT’s output by randomly sampling 18 generated test cases across different configurations. Each case included the modified input, ChatGPT’s sentiment prediction, and its explanation. Here is an example of generated input with the labels provided by SUT and ChatGPT’s validation.
Generated Text:
"Unreal and contrived melodrama, with a screenplay altered and altered
by Mel from his play, The Man Business Organization. A room put up and
improving <unk>, taking atomic and atomic amount il, a troubled and
trouble oneself <unk>, who may be homicidal and murderous...",
Expected Label: "Negative",
Predicted Label: "Positive",
ChatGPT’s Answers: "negative. The review expresses a strong sense of
dissatisfaction with the film, using words like ’unreal’, ’contrived’...
These terms indicate a negative sentiment towards the movie, suggesting
that the reviewer found it lacking in quality and engaging elements."
5 human annotators independently assessed whether (i) the predicted sentiment aligned with the input and (ii) the explanation was logically consistent and specific. To quantify the overall agreement among human participants and ChatGPT, we computed Fleiss’ kappa, obtaining a value of 0.71, indicating a substantial inter-rater agreement Fleiss:1971; Landis:1977. In addition to the group-level measure, we assessed the individual consistency of each participant by comparing their responses with those of the others on every task. The results reveal that the assessment of ChatGPT was the most reliable (0.88) with the highest agreement rate, closely followed by other human annotators (0.78-0.88).